# DiaSynth - Synthetic Dialogue Generation Framework

### v1

@@ Abstract
The scarcity of domain-specific dialogue datasets across various domains, from academic topics to everyday conversations, limits the development of dialogue systems for various applications. Existing research is often constrained either by dialogue datasets that are too general or by niche domain dialogue datasets whose scale does not match the required scale for training dialogue systems. To address this gap, we introduce DiaSynth - a synthetic dialogue generation framework capable of generating high-quality, contextually rich dialogues across a wide range of domains. Our approach differs from existing frameworks by dynamically generating dialogues that incorporate simulated personas, subtopics, and diverse conversational characteristics, using a Large Language Model (LLM) with Chain of Thought (CoT) reasoning to create contextually rich, domain-specific dialogues that closely mimic natural human interactions. DiaSynth produces tailored dialogues that emulate realistic conversations. We perform our experiments by generating synthetic data using different LLMs and few-shot examples from DialogSum and SAMSum. The pretrained language models fine-tuned on the synthetic data outperform the base models by 16.47%, while the comparison between models fine-tuned on in-domain data and synthetic data shows that the synthetic data is able to capture 90.48% of the distribution of the in-domain data. The quality of the data generated also scales with the size of LLMs. These results validate DiaSynth’s potential as a robust alternative to traditional data collection methods.
@@ 1 Introduction
The development of dialogue systems is an important area of research in the field of natural language processing, due to its wide range of applications ranging from customer service chatbots to virtual personal assistants. The effectiveness of these systems rely heavily on the availability of high-quality, large-scale, domain-specific dialogue datasets. Unfortunately, the lack of high-quality, large-scale datasets across various domains such as academic discussions, healthcare consultations and everyday conversations, poses a big challenge. This scarcity limits the ability of developing dialogue systems that can generalize well across diverse settings and domains.
@@
Existing research has attempted to address this issue by collecting domain specific and real life conversations but they fall into one of the two problematic categories. On the one hand, the conversations in a domain specific dataset are superficial and do not go deep into the domain. On the other hand, niche domain dialogue datasets, while contextually rich, often suffer from limited scale. This imbalance does not allow us to develop dialogue systems for underrepresented domains, where collecting large-scale, high-quality data is either costly or logistically complex.
@@
To address these problems, we introduce DiaSynth, a synthetic dialogue generation framework that produces contextually rich and realistic dialogues tailored to specific domains. DiaSynth, with a Large Language Model (LLM) as the backbone, produces high-quality conversations by simulating personas (representing the characteristics and perspectives of a person) and characteristics, which emulates the implicit qualities of a conversation, such as tone, formality, and agreement levels. By using an LLM as the backbone and Chain of Thoughts (CoT) Wei et al. (2024) as the reasoning mechanism, DiaSynth dynamically generates dialogues that closely mimic real-world conversations, allowing for the rapid generation of tailored dialogue data across a wide range of domains. This approach not only addresses the issue of lack of domain specific dialogue data but also provides a scalable, cost-effective alternative to traditional data collection methods.
@@
To validate the effectiveness of DiaSynth, we evaluated the framework on two criteria: the quality of the generated data and the usability of this data for downstream tasks. The results for the quality criterion showed that the data quality improved with the scale of the model. For the usability criterion, we tested the framework on dialogue summarization task. The results demonstrate that models fine-tuned on DiaSynth-generated data outperformed their base versions by an average of 16.47% across multiple evaluation metrics. Additionally, our comparative analysis reveals that synthetic data captures approximately 90.48% of the performance achieved when fine-tuning on in-domain data, underscoring the framework’s potential as a strong alternate when domain-specific data is unavailable.
@@
The remainder of this paper is organized as follows: Section 2 reviews related work on dialogue datasets and synthetic data generation techniques. Section 3 details the DiaSynth framework, including its architecture and data generation methodology. Section 4 presents our experimental setup, datasets, and evaluation metrics. Section 5 discusses the results of our experiments, comparing the performance of models fine-tuned on DiaSynth-generated data versus in-domain data. Section 6 discusses the limitations of DiaSynth and potential future directions for this research. Finally, Section 7 concludes with a summary of our findings.
@@ 3 DiaSynth
DiaSynth is a synthetic dialogue generation framework designed to address the scarcity of high-quality, large-scale, domain-specific dialogue datasets. The framework functions with an LLM as the backbone and CoT as the reasoning mechanism, to simulate diverse real world dialogues that capture the complexity and nuances of human conversations.
DiaSynth operates by taking in a list of topics from the user based on which the conversations should be generated. The users can optionally provide few-shot examples of the format in which they want the dialogue to be generated. Dialogues that are generated directly from the given list of topics will be too superficial since the topics given by the user will not be specific. To overcome this lack of specificity, we generate $m$ sub topics for each of the $n$ topics given by the user. Generating dialogues from the subtopics will have specificity but the dialogues will lack variety. This is because every dialogue is influenced implicitly by the personas of the people involved in the dialogue and, other characteristics such as the location, emotion and more. To increase both variety and depth in the dialogues generated, we generate $p$ personas for each of the $m$ subtopics. The dialogues are then generated for all possible combinations of personas belonging to each of the subtopics. To further ground the dialogues in various settings and characteristics, we employ CoT reasoning during the generation process. DiaSynth employs CoT to reason about the settings and characteristics of a dialogue, as listed in Table 1, ensuring that the dialogues are contextually rich and realistic. This three-stage pipeline not only guarantees the quality of the generated dialogues but also allows for exponential scalability, as illustrated programmatically in Appendix A.
3.1 Subtopic Generation
3.2 Persona Generation
Personas of the individuals involved in a conversation are primary influencers in determining how a conversation pans out. Using random personas from persona datasets and prompting the LLM to simulate a dialogue between them about a random topic often leads to superficial dialogues that lack depth and contextual richness. To address this issue, DiaSynth generates a user-specified number of personas for each subtopic, ensuring that the personas are conditioned on the subtopic context. This conditioning prompts the LLM to create personas that are most likely to engage in a meaningful dialogue about the subtopic, such as a medical professional and a patient discussing "medical diagnostics" or a researcher and a student talking about "academic publishing." We also run a similarity check for the personas too. The conditioned persona generation is crucial because it ensures that future dialogues will not only be contextually rich but also exhibit a high level of depth. Each dialogue will be between two personas who have relevant expertise or perspectives on the given subtopic, allowing the conversation to explore nuances that would otherwise be missed in a generic dialogue setting. Notably, during our experiments, the similarity checks indicated no significant overlap among the dialogues, suggesting that the inherent diversity of the dialogues stems from DiaSynth’s multi-stage generation pipeline.
Figure 1: Overview of the Dialogue Synthesis Framework: The diagram shows the step-by-step process of generating dialogues from user-provided topics, involving sub-topic generation, persona generation, dialogue generation, and filtering based on calculated metrics for downstream tasks.
3.3 Dialogue Generation
The final stage in DiaSynth’s pipeline is the generation of dialogues, where all the components—subtopics, personas, and characteristics—converge to create contextually rich and realistic conversations. This step uses an LLM as the backbone and CoT as the reasoning mechanism, allowing the model to simulate dialogues that incorporate various aspects of human interaction. In this stage, for each subtopic, DiaSynth generates dialogues by pairing every possible combination of personas within that subtopic. The generation process also takes into account the predefined dialogue characteristics as listed in Table 1—such as emotional state, formality level, and familiarity between speakers—to guide the conversation flow and style. These characteristics are explicitly defined in the CoT prompt, allowing the LLM to generate dialogues that are both contextually appropriate and realistic.
Characteristic Description
Age and Gender Defines the demographic details of both personas, influencing the style, tone, and cultural references in the dialogue.
Familiarity Level Indicates how well the speakers know each other, which affects the formality, trust, and depth of the conversation.
Emotional States Specifies the emotional condition of the speakers (e.g., angry, happy, sad), impacting the tone and progression of the dialogue.
Formality Level Determines the level of politeness or casualness in the dialogue, which depends on context and familiarity between speakers.
Duration of the Conversation Suggests the intended length of the conversation, which can influence the dialogue’s depth and complexity.
Communication Medium Defines the medium of the conversation (e.g., face-to-face, phone, text), which affects the dialogue style and expressions used.
Topic of the Conversation Provides the main subject around which the dialogue is centered, guiding the content and direction of the conversation.
Location of the Conversation Indicates where the dialogue takes place, adding context that can influence the dialogue’s formality and content.
Agreement or Disagreement Shows whether the speakers agree or disagree on the topic, which drives the dialogue dynamics and interaction style.
Natural Dialogue Features Incorporates elements like fillers (e.g., "umm," "uh"), pauses, and slang, adding authenticity and naturalness to the dialogue.
Table 1: Characteristics of the Dialogue for CoT Prompt
4 Experimental Setup
In this section, we detail the experimental setup used to evaluate the effectiveness of DiaSynth. Our evaluation focuses on two criteria - quality of the dialogues generated and usability of the dialogues generated for a downstream task. Quality of the dialogues is evaluated using metrics such as FED, GPTScore, and G-Eval. The usability of the dialogues generated for a downstream task is evaluated by choosing summarization as the downstream task.
4.1 Quality of the dialogues
To evaluate the quality of the dialogues, we employ the metrics that have been developed for evaluating the quality of text generated by LLMs. The metrics are as follows,
  • FED Mehri and Eskenazi (2020) - FED evaluates the quality of dialogue by appending a positive and a negative utterance to the end of the dialogue and measuring the difference in probabilities assigned to the positive and negative utterances.
  • GPTScore Fu et al. (2024) - GPTScore assesses dialogue quality based on various characteristics such as coherence, diversity, and others. The evaluation involves asking an LLM whether a dialogue meets a specific criterion. The score for each criterion is determined by the probability of the LLM responding affirmatively (’Yes’).
  • G-Eval Liu et al. (2023) - G-Eval evaluates dialogue quality by prompting an LLM to rate the dialogue on a scale of 1 to 3 across multiple criteria. The final score is calculated using a weighted average of each rating, with weights derived from the probability distribution of the LLM’s assigned ratings.
To validate the framework across models and also domains, we generate data using three open source LLMs, one closed source LLM and also use few shot examples from two different dialogue datasets. The four LLMs used are - Phi-3 (phi-3-mini-128k-instruct (3.8B)  [^1] Phi-3 HF link) Abdin et al. (2024), InternLM-2.5 (internlm2_5-7b-chat (7.74B) [^2] InternLM-2.5 HF link) Cai et al. (2024), LLaMA-3 (llama3-8b-instruct (8.03B) [^3] LLaMA-3 HF link) Dubey et al. (2024) and GPT-4o. The two different dialogue datasets that were used are DialogueSum Chen et al. (2021) and SAMSum Gliwa et al. (2019)
4.2 Downstream Task - Summarization
To evaluate the usability of the dialogues generated by DiaSynth, we choose summarization as the downstream task. Summarization is an essential application of dialogue systems, where the goal is to generate concise, and contextually relevant summaries. We use three established evaluation metrics—QAGS, BERTScore, and ROUGE-L—to assess the performance of summarization models fine-tuned on DiaSynth-generated data.
  • QAGS Wang et al. (2020) - QAGS (Question Answering Guided Summarization) evaluates the factual consistency of generated summaries by automatically generating questions based on the summary and then comparing the answers obtained from the summary to those from the source dialogue. In our work, instead of generating questions from the summaries, we generate them from the source dialogues, allowing us to directly compare the consistency and accuracy of summaries generated from different datasets containing the same dialogues.
  • BERTScore Zhang* et al. (2020) - BERTScore measures the semantic similarity between the generated summary and the reference summary using contextual embeddings.
  • ROUGE-L Lin (2004) - ROUGE-L is a widely used evaluation metric that measures the overlap of longest common subsequences (LCS) between the generated summary and the reference summary.
We fine-tune pretrained models that perform well on summarization such as DistilBART, BART Lewis et al. (2020), T5 Raffel et al. (2020) and LED Beltagy et al. (2020), on DiaSynth-generated dialogues and evaluate their performance using the above metrics. We evaluate the effectiveness of DiaSynth in two key aspects: first, by assessing the performance improvement of models fine-tuned on DiaSynth-generated data compared to the pretrained models; and second, by measuring the extent to which DiaSynth-generated data reflects real-world data distribution by comparing the performance of models fine-tuned on DiaSynth data versus those fine-tuned on in-domain data.
5 Results
This section discusses the results of the data generated using DiaSynth (quality of the data and usability in downstream tasks) with different LLMs as the backbone and varying few-shot examples. Specifically, we utilized Phi-3, InternLM-2.5, LLaMA-3 and GPT-4o as the LLM backbones, and the few-shot examples were sourced from DialogueSum and SAMSum datasets. These combinations allow us to evaluate the robustness and adaptability of DiaSynth across different models and few shot examples. In total, eight distinct datasets were generated using DiaSynth by pairing each LLM backbone with the two sets of few-shot examples, resulting in all possible combinations. For each dataset, DiaSynth was provided with the same 16 broad topics and tasked with generating 6 subtopics for each topic, followed by creating 6 personas for each subtopic. The statistics of the datasets generated using DiaSynth, including the number of dialogues, average number of turns, and average number of tokens per turn, are summarized in Table 2.
LLM Few-Shot Number of Samples Avg. number of turns Avg. number of tokens per turn Diversity (rouge-l)
Phi-3 DialogueSum 1215 9.13 20.38 0.27
Phi-3 SAMSum 1410 13.98 13.94 0.27
InternLM-2.5 DialogueSum 1035 9.23 27.98 0.30
InternLM-2.5 SAMSum 1135 13.96 19.07 0.29
LLaMA-3 DialogueSum 1154 6.86 31.99 0.29
LLaMA-3 SAMSum 1195 10.54 20.41 0.29
GPT-4o DialogSum 1375 15.16 15.96 0.29
GPT-4o SAMSum 1380 15.43 13.53 to 0.28
Table 2: Data Statistics
5.1 Quality of the Dialogues
The quality of the synthetic datasets produced by DiaSynth was evaluated using FED, GPTScore, and G-Eval metrics, as detailed in Tables 3, 4, and 5. The results reveal distinct variations in performance across different model and dataset configurations, reflecting the unique characteristics of each.
The FED scores in Table 3 indicate that LLaMA-3 and GPT-4o consistently outperform Phi-3 and InternLM-2.5, particularly in coherence, consistency, and diversity, with LLaMA-3 leading in the DialogSum dataset (e.g., coherence 0.9684, consistency 0.9570, diversity 0.9596) and GPT-4o excelling in the SAMSum dataset, especially in coherence (0.9939) and consistency (0.9878). Table 4 further supports LLaMA-3’s adaptability and error-handling abilities, as it outperforms others in diversity (0.0430) and inquisitiveness (0.0415) in DialogSum, while maintaining an edge in SAMSum for diversity (0.0655) and inquisitiveness (0.0622). Though GPT-4o remains strong in coherence and consistency, Phi-3 shows competitive coherence in SAMSum (0.0325). Finally, Table 5 highlights GPT-4o’s dominance in engagingness and naturalness with perfect scores (3.0) for DialogSum, while InternLM-2.5 stands out in coherence (2.9990) and groundedness (2.9973) for DialogSum, and coherence (2.9983) and groundedness (2.9952) for SAMSum, suggesting it maintains high factual accuracy. The contrasting performance of GPT-4o on the DialogSum and SAMSum datasets can be attributed to the differing structures of the dialogues in these datasets. DialogSum consists of more formal and structured dialogues, which aligns with the typical response style of GPT-4o, leading to its stronger performance. In contrast, SAMSum contains more casual, human-like conversations, which might explain GPT-4o’s relatively poorer performance, as it may not adapt as well to the informal, spontaneous nature of such dialogues. Overall, while GPT-4o excels in natural and engaging dialogue, LLaMA-3 offers the most versatility, and InternLM-2.5 provides a strong alternative with high coherence and groundedness.

### v2

# DiaSynth: Synthetic Dialogue Generation Framework for Low Resource Dialogue Applications

@@ Abstract
The scarcity of domain-specific dialogue datasets limits the development of dialogue systems across applications. Existing research is constrained by general or niche datasets that lack sufficient scale for training dialogue systems. To address this gap, we introduce DiaSynth - a synthetic dialogue generation framework capable of generating high-quality, contextually rich dialogues across a wide range of domains. Unlike existing frameworks, DiaSynth uses Large Language Models (LLMs) and Chain of Thought (CoT) reasoning to generate dynamic, domain-specific dialogues with simulated personas and diverse conversational features. We perform our experiments by generating synthetic data using different LLMs and few-shot examples from DialogSum and SAMSum. The pretrained language models fine-tuned on the synthetic data outperform the base models by 16.47% on dialogue summarization, while the comparison between models fine-tuned on in-domain data and synthetic data shows that the synthetic data is able to capture 90.48% of the performance distribution of the in-domain data on dialogue summarization. The quality of the data generated also increases as we increase the size of LLM from 3B to 8B. These results validate DiaSynth’s potential as a robust alternative to traditional data collection methods. We open source the code and data generated for future research.
@@ 1 Introduction
Dialogue systems are crucial in natural language processing due to applications like customer service chatbots and virtual assistants. Their effectiveness depends on large, high-quality, domain-specific datasets. The lack of large-scale, high-quality datasets across domains like academic discussions, healthcare, and everyday conversations poses a challenge. This scarcity limits the development of dialogue systems that generalize well across domains.
@@
Prior work (Feng et al. (2020), Zeng et al. (2020), Budzianowski et al. (2018)) collects domain-specific dialogues but often lacks depth, scale, or domain diversity. On the one hand, the conversations in a domain specific dataset are superficial and do not go deep into the domain. On the other hand, niche domain dialogue datasets, while contextually rich, often suffer from limited scale. This imbalance hinders dialogue system development in underrepresented domains, where data collection is costly and complex.
@@
To address these problems, we introduce DiaSynth, a synthetic dialogue generation framework that produces contextually rich and realistic dialogues tailored to specific domains. DiaSynth, using a Large Language Model (LLM), generates high-quality conversations by simulating personas and conversation characteristics like tone and formality. With LLMs and Chain of Thoughts (CoT) Wei et al. (2024), DiaSynth generates dialogues that mimic real-world conversations for a wide range of domains. CoT plays a crucial part by building different environments for different personas (refer to Appendix D for examples) which influence varied conversations. This approach addresses data scarcity and offers a scalable, cost-effective alternative to traditional methods.
@@
To validate the effectiveness of DiaSynth, we evaluated the framework on two criteria: the quality of the generated data and the usability of this data for downstream tasks. The results for the quality criterion showed that the data quality improved with the scale of the model. For usability, we tested DiaSynth on dialogue summarization. Models fine-tuned on DiaSynth data outperformed base versions by 16.47% on average. Additionally, DiaSynth’s synthetic data captures 90.48% of in-domain performance, highlighting its potential as a strong alternative when domain-specific data is unavailable.
@@
The remainder of this paper is organized as follows: Section 2 reviews related work on dialogue datasets and synthetic data. Section 3 details the DiaSynth framework and methodology. Section 4 describes our experimental setup and evaluation. Section 5 compares the performance of models fine-tuned on DiaSynth data to in-domain data. Section 6 concludes with a summary of our findings and Section 7 discusses the limitations of DiaSynth and potential future directions for this research.
@@ 2.2 Prompting Task-Oriented Dialogue Generation
LLM Few-shot examples Number of Samples Avg. number of turns Avg. number of tokens per turn Diversity (ROUGE-L)
Phi-3 DialogSum 1215 9.13 20.38 0.27
InternLM-2.5 DialogSum 1035 9.23 27.98 0.30
LLaMA-3 DialogSum 1154 6.86 31.99 0.29
GPT-4o DialogSum 1375 15.16 15.96 0.29
Phi-3 SAMSum 1410 13.98 13.94 0.27
InternLM-2.5 SAMSum 1135 13.96 19.07 0.29
LLaMA-3 SAMSum 1195 10.54 20.41 0.29
GPT-4o SAMSum 1380 15.43 13.53 0.28
Table 1: Data Statistics
@@
2.3 Existing Task-Oriented Dialogue Datasets
In addition to prompt-based synthetic data generation, various large-scale dialogue datasets have been instrumental in advancing task-oriented dialogue systems. Among these, the MultiWOZ dataset Budzianowski et al. (2018) is a prominent resource, providing richly annotated dialogues across multiple domains. MultiWOZ has enabled researchers to train models capable of handling complex, multi-turn interactions across diverse tasks. The nature of MultiWOZ’s annotations has made it a benchmark for evaluating the performance of dialogue systems, though it is often complemented by synthetic data to introduce further diversity and variation in dialogue scenarios.
Similarly, Doc2Dial Feng et al. (2020) is another widely used dataset designed specifically for document-grounded dialogue systems. Doc2Dial includes conversations grounded in structured documents, focusing on providing users with accurate and relevant information based on their inquiries. This dataset has been instrumental in improving the ability of dialogue systems to retrieve and generate accurate responses when interacting with complex information sources. However, much like MultiWOZ, Doc2Dial’s scope is limited to the predefined topics and domains covered within the dataset, which can restrict model generalizability to new or unseen situations.
To overcome the domain-specific limitations of datasets, our study adopts a synthetic data generation approach that expands existing topics into subtopics, thus providing a broader and deeper pool of conversational data. By incorporating both task-oriented prompts and personas, our generated dialogues aim to complement these datasets by offering more personalized and contextually rich conversations, thereby enhancing the robustness and versatility of dialogue systems
@@ 3 DiaSynth
DiaSynth is a synthetic dialogue generation framework designed to address the scarcity of high-quality, large-scale, domain-specific dialogue datasets. DiaSynth uses an LLM and CoT reasoning to simulate diverse, nuanced dialogues.
DiaSynth takes a list of user-provided topics to generate dialogues. The users can optionally provide few-shot examples of the format in which they want the dialogue to be generated. Directly generating dialogues from user topics would be too superficial due to their lack of specificity. To overcome this lack of specificity, we generate $m$ sub topics for each of the $n$ topics given by the user. Generating dialogues from the subtopics will have specificity but the dialogues will lack variety. This is because every dialogue is influenced implicitly by the personas of the people involved in the dialogue and, other characteristics such as the location, emotion and more. To enhance variety and depth, we generate $p$ personas per subtopic and create dialogues for all persona-subtopic combinations. To further ground the dialogues in various settings and characteristics, we employ CoT reasoning during the generation process. DiaSynth employs CoT to reason about the settings and characteristics of a dialogue, which are listed in Appendix C, ensuring that the dialogues are contextually rich and realistic. This three-stage pipeline not only guarantees the quality of the generated dialogues but also allows for exponential scalability, as illustrated programmatically in Appendix B.
3.1 Subtopic Generation
3.2 Persona Generation
Personas of the individuals involved in a conversation are primary influencers in determining how a conversation pans out. Using random personas from persona datasets and prompting the LLM to simulate a dialogue between them about a random topic often leads to superficial dialogues that lack depth and contextual richness. To address this issue, DiaSynth generates a user-specified number of personas for each subtopic, ensuring that the personas are conditioned on the subtopic context. This conditioning prompts the LLM to create personas that are most likely to engage in a meaningful dialogue about the subtopic, such as a medical professional and a patient discussing "medical diagnostics" or a researcher and a student talking about "academic publishing." We also run a similarity check for the personas too. The conditioned persona generation is crucial because it ensures that future dialogues will not only be contextually rich but also exhibit a high level of depth. Each dialogue will be between two personas who have relevant expertise or perspectives on the given subtopic, allowing the conversation to explore nuances that would otherwise be missed in a generic dialogue setting.
3.3 Dialogue Generation
The final stage in DiaSynth’s pipeline is the generation of dialogues, where all the components—subtopics, personas, and characteristics—converge to create contextually rich and realistic conversations. This step uses an LLM as the backbone and CoT as the reasoning mechanism, allowing the model to simulate dialogues that incorporate various aspects of human interaction. DiaSynth generates dialogues by pairing all persona-subtopic combinations. The process also integrates predefined characteristics (Table 11) like emotional state, formality, and familiarity to guide the flow and style. These characteristics are defined in the CoT prompt, guiding the LLM to generate realistic, contextually appropriate dialogues. The importance of CoT and the lack of it affects the quality of the dialogues, which is shown quantitatively in Appendix E.
4 Experimental Setup
@@
In this section, we detail the experimental setup used to evaluate the effectiveness of DiaSynth. Our evaluation focuses on two criteria - quality of the dialogues generated and usability of the dialogues generated for a downstream task. Quality of the dialogues is evaluated using metrics such as FED, GPTScore, and G-Eval. We evaluate the usability of DiaSynth-generated dialogues by using summarization as the downstream task.
4.1 Quality of the dialogues
To evaluate the quality of the dialogues, we employ the metrics that have been developed for evaluating the quality of text generated by LLMs. We use the following metrics:
  • FED Mehri and Eskenazi (2020) - FED evaluates dialogue quality by comparing the probabilities of appended positive and negative utterances.
  • GPTScore Fu et al. (2024) - GPTScore assesses dialogue quality by asking an LLM to evaluate criteria like coherence and diversity, with scores based on the probability of affirmative responses.
  • G-Eval Liu et al. (2023) - G-Eval rates dialogue on a 1-3 scale across criteria, with final scores as a weighted average from the LLM’s probability distribution.
To validate the framework across models and also domains, we generate data using three open source LLMs, one closed source LLM and also use few shot examples from two different dialogue datasets. The open sourced LLMs are - Phi-3 Abdin et al. (2024), InternLM-2.5 Cai et al. (2024), LLaMA-3 Dubey et al. (2024) and the closed source LLM used is GPT-4o. The 8-bit quantized versions of the open source LLMs were used for faster experimentation and generation. The two different dialogue datasets that were used as few-shot examples are DialogSum Chen et al. (2021) and SAMSum Gliwa et al. (2019)
4.2 Downstream Task - Summarization
To evaluate the usability of the dialogues generated by DiaSynth, we choose summarization as the downstream task. Summarization, a key application of dialogue systems, aims to generate concise, contextually relevant summaries. We use three established evaluation metrics—QAGS, BERTScore, and ROUGE-L—to assess the performance of summarization models fine-tuned on DiaSynth-generated data.
  • QAGS Wang et al. (2020) - QAGS evaluates factual consistency by generating questions from the summary and comparing answers to those from the source dialogue.
  • BERTScore Zhang* et al. (2020) - BERTScore measures semantic similarity between generated and reference summaries using contextual embeddings.
  • ROUGE-L Lin (2004) - ROUGE-L measures longest common subsequence (LCS) overlap between generated and reference summaries.
We fine-tune pretrained summarization models like DistilBART, BART Lewis et al. (2020), T5 Raffel et al. (2020) and LED Beltagy et al. (2020), on DiaSynth-generated dialogues and evaluate their performance using the above metrics. We evaluate the usability of DiaSynth in two key aspects: first, by assessing the performance improvement of models fine-tuned on DiaSynth-generated data compared to the pretrained models; and second, by measuring the extent to which DiaSynth-generated data reflects real-world data distribution by comparing the performance of models fine-tuned on DiaSynth data versus those fine-tuned on in-domain data.
5 Results
This section discusses the results of the data generated using DiaSynth (quality of the data and usability in downstream tasks) with different LLMs and varying few-shot examples. Specifically, we utilized Phi-3, InternLM-2.5, LLaMA-3 and GPT-4o as the LLM backbones, and the few-shot examples were sourced from DialogueSum and SAMSum datasets. These combinations allow us to evaluate the robustness and adaptability of DiaSynth across different models and few shot examples. In total, eight distinct datasets were generated using DiaSynth by pairing each LLM with the two sets of few-shot examples, resulting in all possible combinations. For each combination, DiaSynth was provided with the same 16 broad topics and tasked with generating 6 subtopics for each topic, followed by creating 6 personas for each subtopic. The statistics of the datasets generated using DiaSynth, including the number of dialogues, average number of turns, and average number of tokens per turn, are summarized in Table 1. All the experiments were run on a single A100 GPU with the generation time ranging from 2 hours to 4 hours.
5.1 Quality of the Dialogues
The quality of the synthetic datasets produced by DiaSynth was evaluated using FED, GPTScore, and G-Eval metrics, as detailed in Tables 2, 3, and 4. The results reveal distinct variations in performance across different model and dataset configurations, reflecting the unique characteristics of each.
5.1.1 Metric Scores
FED: The FED scores in Table 2 show that LLaMA-3 and GPT-4o achieve almost a perfect score (+1) in most of the criteria, while Phi-3 and InternLM-2.5 also have decent performances. GPT-4o has a clear advantage when it comes to generating likeable dialogues while there is not much separation on other criteria.
GPTScore: Results illustrated in 3 are surprising in that GPT-4o is the worst performing model on GPTScore, which might require further research while LLaMA-3 clearly dominates the other models.
G-Eval: Table 4 highlights GPT-4o’s dominance in engagingness and naturalness with perfect scores (3.0) for DialogSum, while InternLM-2.5 stands out in coherence (2.9990) and groundedness (2.9973) for DialogSum, and coherence (2.9983) and groundedness (2.9952) for SAMSum, suggesting it maintains high factual accuracy.
Dataset-Specific Performance. The contrasting performance of GPT-4o on the DialogSum and SAMSum datasets in Table 2 can be attributed to the differing structures of the dialogues in these datasets. DialogSum consists of more formal and structured dialogues, which aligns with the typical response style of GPT-4o, leading to its stronger performance. In contrast, SAMSum contains more casual, human-like conversations, which might explain GPT-4o’s relatively poorer performance, as it may not adapt as well to the informal, spontaneous nature of such dialogues. Overall, while GPT-4o excels in natural and engaging dialogue, LLaMA-3 offers the most versatility, and InternLM-2.5 provides a strong alternative with high coherence and groundedness.